# Adaptive Pruning for Increased Robustness and Reduced Computational Overhead in Gaussian Process Accelerated Saddle Point Searches

### v1

@@ 1 Introduction
Several methods have been developed for finding first order saddle points on energy surfaces given only a starting configuration of the atoms and without a predefined final state of the transition. [7, 54, 38, 20, 24, 39, 48, 22] A given initial state, without reference to a known product state, can in general lead to many different saddle points. By making a catalog of the lower energy saddle points, which correspond to the faster transitions, long time scale simulations of the system can be carried out for a given temperature. [21, 5] But, the computational effort involved when the PES is obtained from an electronic structure calculation makes this intractable. General purpose machine learned interatomic potentials, such as PET-MAD, [35, 6] can alleviate some of this cost for systems where a large set of electronic structure calculations has been performed, but these often suffer from under-sampling of the regions near the transition states as they are of low probability and are harder to sample.[50].
@@ 2.2 Optimal Transport GP
The robustness and computational scaling of GP-accelerated searches link directly to the metric used to define dissimilarity between molecular configurations. To address these challenges, we introduce the Optimal Transport Gaussian Process (OT-GP) framework, which employs a metric grounded in Optimal Transport (OT) theory. OT provides a powerful means to “lift” a simple geometric distance between points to a more sophisticated distance between entire distributions. We treat each molecular configuration as a discrete probability distribution of element-typed (“colored”) points in three-dimensional space. The distance between two such configurations then corresponds to the minimum “work” required to transport one arrangement of atoms to the other. This concept, formally the Wasserstein-1 distance, finds its computational analogue in the Earth Mover’s Distance (EMD) [52, 29]. Figure 2 summarizes the key modifications to the algorithm workflow in this improved GP acceleration of saddle point searches.
@@ 4.3 Hyperparameter Guardrails
Embedding the barrier of Eq (16}) directly into the objective prevents $\sigma_{f}^{2}$ from exceeding $\exp(\lambda_{\max})$. Because the barrier strength $\mu$ scales with the data count, Eq. (17}), the surrogate remains flexible in the early steps. This allows the optimisation to explore a wide variance range and progressively tightens as the data set matures, exactly when the model should have settled on a physically reasonable amplitude. This adaptive behaviour eliminates the pathological growth of the variance while preserving the surrogate model’s ability to capture the curvature of the true PES, as seen in Figure 4 (B).
@@ Acknowledgments
S5.2 Post-Initialization Comparability for D136
@@
S6.1 Model specifications
S6.1.1 Prior distributions
@@
S6.1.4 Total wall-clock time

### v2

@@ 1 Introduction
Several methods have been developed for finding first order saddle points on energy surfaces given only a starting configuration of the atoms [8, 59, 43, 23, 28, 44, 53, 26] without a predefined final state of the transition. These are distinct from the “double ended” methods like the Nudged elastic band [29, 48, 25, 17] and variants thereof [2] which require known initial and final states. We focus on general-purpose methods which do not require user-defined constraints on the transition state optimization [14] and compare to the extant state of the art within such methods [19]. A given initial state, without reference to a known product state, can in general lead to many different saddle points. By making a catalog of the lower energy saddle points, which correspond to the faster transitions, long time scale simulations of the system can be carried out for a given temperature. [24, 6] But, the computational effort involved when the PES is obtained from an electronic structure calculation makes this intractable. General purpose machine learned interatomic potentials, such as PET-MAD, [40, 7] can alleviate some of this cost for systems where a large set of electronic structure calculations has been performed, but these often suffer from under-sampling of the regions near the transition states as they are of low probability and are harder to sample.[55].
@@ 2.2 Optimal Transport GP
The robustness and computational scaling of GP-accelerated searches link directly to the metric used to define dissimilarity between molecular configurations, $D(\mathbf{x}_{i},\mathbf{x}_{j})$. Metrics that depend on fixed atom indices, such as the “1D max log” metric (Eq. 7), lack invariance to the permutation of identical atoms. Even low-energy events, like rotations of a methyl group, may swap indices of two hydrogen atoms, resulting in a large, unphysical distance in this metric. To solve this, we ground our dissimilarity measure in Optimal Transport (OT) theory [57, 34]. OT provides a deterministic prescription to “lift” a geometric distance between points into a distance between entire distributions of points. We treat each molecular configuration $\mathbf{x}$ as a discrete distribution of element-typed (i.e., “colored”) points in three-dimensional space. Conceptually, OT frames the distance calculation as a “transport problem.” If we consider a configuration $\mathbf{x}_{i}$ to be piles of “mass” (the atoms) and configuration $\mathbf{x}_{j}$ as a set of “holes” into which the mass must be moved, then the cost of moving one unit of mass from a point $\mathbf{r}_{a}$ to a point $\mathbf{r}_{b}$ is the Euclidean distance $\|\mathbf{r}_{a}-\mathbf{r}_{b}\|$. The Earth Mover’s Distance (EMD), the computational analogue of the Wasserstein-1 distance, finds the “transport plan” that minimizes the total work, i.e., the (mass $\times$ distance) summed over all movements. By enforcing coloring, we have that “carbon” masses only move to “carbon” holes, respecting chemical identity. More importantly, the EMD inherently solves the permutation problem as the optimal transport plan corresponds to the optimal matching between the two sets of atoms, automatically assigning atom 1 in $\mathbf{x}_{i}$ to its “correct” counterpart in $\mathbf{x}_{j}$, regardless of its index. This yields a physically meaningful distance that correctly identifies symmetrically equivalent configurations as identical. We leverage this robust, permutation-invariant metric in our Farthest Point Sampling (FPS) to select a truly diverse set of configurations for hyperparameter optimization. Figure 2 summarizes the key modifications to the algorithm workflow in this improved GP acceleration of saddle point searches.
@@ 3.1.1 Reliability
We further investigate the four systems (Fig. 3 (A), blue dots) where the GPDimer runs succeed but the OTGPD runs fail. This discrepancy does not stem from a flaw in the OTGPD algorithm but from an artifact of the dimer initialization procedure (visualized in Fig. 2). For OTGPD runs, we form the initial dimer configuration by displacing atoms along the softest mode found by an inexpensive semi-empirical (xtb, GFN2-xTB) calculation. In these specific cases, this displacement produced an unphysical starting geometry with high atomic forces or atoms in close contact. When any GP-accelerated method begins from such a high-energy, high-force baseline, the initial surrogate model learns high forces and energies corresponding to unphysical geometries are not anomalous, thus “poisoning” the run, leading to optimization failures. The GPDimer benchmark used a different initialization method which, for these systems, happened to produce a viable starting configuration. Robust dimer initialization routines are not a solved problem, nor the focus of this work. As detailed in the SI, when the OTGPD method starts from the same initial geometries used by the GPDimer, it converges successfully and more efficiently. Across these four systems, the re-run OTGPD required a near-identical total number of HF evaluations (146 for OTGPD vs. 148 for GPDimer) while completing in less than half the average wall time (OTGPD, 51.3 minutes vs. GPDimer, 109.8 minutes), despite running with half the number of parallel processes (OTGPD 8 vs GPDimer 16).
@@ 4.1 Improved Wall-time Efficiency
The most common approach to handling GPR scaling involves sparse (inducing point) methods [52, 20]. These methods work well for large, static datasets, however, our “on-the-fly” design lies within a fundamentally different, dynamic regime (Fig. 2). The dynamic re-optimization reduces the advantages of standard sparse GPR which involves an additional optimization at each training step to construct the inducing point locations. We estimated this would prove prohibitive and likely negate any savings, given our median sample size of approximately 30 PES evaluations.
The FPS method confines the expensive hyperparameter optimization to a small subset ($M_{\text{sub}}$) while still using the full data history ($\mathcal{X}$) for the predictive step. This “data selection” approach maintains a link to the physical system as the FPS subset consists of observed configurations that show maximal diversity in EMD metric. These points “bracket” the relevant region of the PES, an interpretation which is harder in a sparse-GP context, where the inducing points represent optimized mathematical constructs. We anticipate sparse methods may prove beneficial for much larger problems requiring hundreds of evaluations, but not for the class of problems considered here.
@@ 4.3 Hyperparameter Guardrails
Embedding the logarithmic barrier of Eq. (16) directly into the objective function serves as a “hard wall” to enforce this constraint. Unlike a soft penalty, the barrier creates a diverging gradient (Eq. 18) as the optimizer approaches the bound $\lambda_{\max}$, which guarantees that the optimizer cannot enter the unstable region. Such constraints on $\sigma_{f}^{2}$ are interpretable, with a physically meaningful upper bound rather than a more abstract L1/L2 penalty strength. We found this gradient-based barrier necessary as a simple hard bound proved insufficient, with the optimizer “sticking” to the boundary.
Because the barrier strength $\mu$ scales with the data count, Eq. (17}), the surrogate remains flexible in the early steps. This allows the optimisation to explore a wide variance range and progressively tightens as the data set matures, exactly when the model should have settled on a physically reasonable amplitude. This adaptive behaviour eliminates the pathological growth of the variance while preserving the surrogate model’s ability to capture the curvature of the true PES, as seen in Figure 4 (B).
@@ Acknowledgments
Github
Materials Archive
Appendix S1 Reproduction note
The full set of benchmark inputs, raw outputs, analysis scripts, and pinned runtime environments used in this study are publicly archived. The GitHub repository and Materials Cloud archive contain the original runs for all 238 systems, the scripts used to generate every figure and table, and environment specifications (container/environment manifests) that reproduce the computational environment.
To diagnose a small number of non-convergent cases we performed local re-runs only for the four systems which fail for the OTGPD but not the GPDimer; these runs and their logs are included in the same archive and flagged in the repository for convenience. Users wishing to reproduce any specific experiment can either use the provided raw outputs, the pre-processed FAIR formatted csv data, or re-run the workflow using the included environment manifests and Snakemake pipelines; exact instructions and file paths are given in the archive and on the Github repository.
@@
S6.2 Post-Initialization Comparability for D136
@@
S7.1 Model specifications
S7.1.1 Prior distributions
@@
S7.1.4 Total wall-clock time
@@
Appendix S8 GPDimer / OTGPD failure modes
The OTGPD framework successfully eliminates the signal variance instability. 4 failures were observed in the OTGPD benchmark but not in the GPDimer which we consider in this section. These were not due to an algorithmic instability but were artifacts of the dimer initialization routine.
For the OTGPD benchmark, initial dimer configurations were generated by displacing atoms along the softest mode found by an inexpensive xtb calculation. For systems D016, D084, D100, and S242, this procedure created a pathological starting geometry with excessively high forces or atoms in unphysically close contact.
When any GP-accelerated method is “poisoned” with such a high-energy, high-force baseline, the initial surrogate model learns that these unphysical configurations are normal, leading to an unstable search that fails to converge. The GPDimer benchmark, used a different (and in these cases, fortuitously non-pathological) initialization based on displacing the least coordinated atom, bypassed this specific problem. Figure S3 shows the initial points at the start of the algorithm runs, extracted from the trajectory files.
Figure S3: Initializations for the GPDimer (top) and OTGPD (bottom). The XTB initialization procedure in these four cases results in starting the runs from unphysical initial geometries leading to failures in the optimization routine. With XTB, D016 has unphysically close carbon atoms, D084 has a shortened carbon-oxygen bond, D100 has near-overlapping carbon atoms, and S242 has a CH3 end-group instead of the hydrogen on the middle carbon.
To confirm this diagnosis, we re-ran these four systems using OTGPD but starting from the exact same (valid) initial configurations used by the GPDimer runs to show all methods converge.
Times reported in minutes –¿ total seconds / 60.
S8.1 D016
Method Calls Time
Dimer 106 7.4
GPDimer 20 3.7
OTGPD 20 3.3
S8.2 D084
Method Calls Time
Dimer 2666 255.4
GPDimer 75 417
OTGPD 65 181.4
S8.3 D100
Method Calls Time
Dimer 214 24.3
GPDimer 28 16.2
OTGPD 28 18.52
S8.4 S242
Method Calls Time
Dimer 249 14.6
GPDimer 25 2.3
OTGPD 33 1.85